Find the equation of the line that passes through the point (4,5) and is parallel to the line y=3x+6.

Answer by robertb(4012)

The line should have the appearance of y = 3x + c, where c is some constant. To solve for c, substitute the coordinates into the equation:
5 = 3*4 + c ==> 5 = 12 + c ==> c = -7.
Therefore the equation is y = 3x - 7.